### v1

# Notes on Correlation Stress Tests

Abstract
This note outlines an approach to stress testing of covariance of financial time series, in the context of financial risk management. It discusses how the geodesic distance between covariance matrices implies a notion of plausibility of covariance stress tests. In this approach, correlation stress tests span a submanifold of constant determinant of the Fisher–Rao manifold of covariance matrices. A parsimonious geometrically invariant definition of arbitrarily large correlation stress tests is proposed, and a few examples are discussed.
1. Introduction
Financial market news sometimes mention changes of correlations between assets in articles with headlines such as: “Correlations Between Credit and Equities Are Breaking Down”[1],“Negative equity/bond correlation is positive for 60-40 portfolio”[2], or “Rates correlations break down amid volatility surge”[3]. Traders and asset managers also refer to correlation breaks, sometimes in a language that implies such breaks are of special interest. In this note we consider several quantitative aspects of changes in correlations and present them in a form suitable for quantitative risk management of financial investments.
What is usually referred to as a “correlation break” is often a simultaneous move of the time series of several markets (in this note also referred to as market factors) that is surprising, given their past covariance. This is likely to be an intuitive assessment of the Mahalanobis distance $d$ from zero of the vector $x$ of the recent market changes with respect to their forecasted covariance matrix $\Sigma$, namely a move which results in a large increase of
$d^{2}=x^{T}\Sigma^{-1}x.$ (1)
However, there is also another notion of a “correlation break”, namely that of realised correlations between markets being meaningfully different from their forecasted correlations.
These two notions are related: if the former type of correlation breaks keep recurring for an extended time, they may then result in the latter. In this note, we will consider the latter type of breaks, and will model such events as a change in the forecasted covariance matrix.
Stress tests are supposed to be unlikely, extreme, but plausible [4, 5, 6]. In our context, “unlikely” probably refers to the low likelihood of a certain covariance matrix given the past data, while “extreme” is bound to mean a large and negative impact on the entity under consideration, in our case a portfolio of financial assets or derivatives. The meaning of plausibility is however less clear. Various approaches to designing and selecting stress tests with reference to the empirical time series of market factors have been proposed. For example, Ref. [7] describes what is perhaps the most common, albeit somewhat arbitrary, practice of stress testing as applied by investment risk management professionals, while Ref. [8] describes stress testing of the correlation matrix of asset returns in a specific parametric form, where correlations are represented as a function of risk factors, such as country and industry factors, and where a sparse factor structure linking assets and risk factors is built using Bayesian variable selection methods.
In the macroeconomic context, Antolin-Diaz et al. [9] developed tools for constructing econometrically meaningful scenarios with structural vector autoregressive models, and proposed a quantitative measure to assess and compare their plausibility, given the macroeconomic data.
In this note we will consider some general aspects that abstract away from the specific interpretation of the underlying data, and concentrate instead of the geometric structure of stress tests. Thus, the methods and measures described here may be applied generically to many problems where stress testing, and particularly correlation stress testing, is of interest. While the approach is generic, we will illustrate the main conclusions with a few example applications.
The Mahalanobis distance provides a useful hint regarding the plausibility of a stress test: perhaps refers to $d$ being relatively large [5]. For the Fisher–Rao manifolds of the multivariate normal probability distributions, the Mahalanobis distance is a special case of the geodesic distance between the distributions: the case when they have different means, but the same covariance. The opposite case, where the multivariate distributions have the same means, but different covariances, will be considered here.
It is intuitively clear what a correlation stress should be: one would like to keep all the variances fixed and change the correlations – ideally just a few or perhaps even just one correlation coefficient – and recombine the so-stressed correlation matrix with the unchanged variances to produce the stressed covariance matrix. This however has a few drawbacks. It can trivially result in a non-positive covariance matrix. Moreover, it may fail to be consistent with no-arbitrage constraints that are present in the markets. However, by far the main drawback of such an approach is its arbitrariness. If one used a different set of factors to describe the exposure of a portfolio, an entirely different correlation stress test would need to be specified. Presumably, the specification of a stress test would be tailored to market factors specified for a portfolio, thus such a correlation stress test would depend implicitly on the portfolio being considered, and also on the somewhat arbitrary selection of the preferred market variables, or market factors, used to define it. Needless to say, it is unlikely that a universal measure of plausibility can be associated with such an arbitrary notion of stress tests.
A change of the risk factors used to describe the portfolio can be thought of as a change of basis of the vector space of market risk factors, at least in the leading order of approximation. Requiring invariance under a change of basis in the vector space of the portfolio risk factors may not appear important at first, but it is essential - a meaningful stress test needs to have the same result when, for example, a fixed-income portfolio that is being stressed is represented using forward rates or alternatively represented using par rates. It is also crucial when some market variables, as a result of no-arbitrage conditions, are related through algebraic identities, as is the case for log-returns of foreign exchange rates. For example, the log-return of the USD/JPY exchange rate added to the log-return of the JPY/GBP exchange rate should be equal to the log-return of the USD/GBP exchange rate, within the accuracy of foreign-exchange arbitrage. Projecting the stress tests on various subspaces of the space of market variables should have consistent results and a measure of the plausibility of a stress test should not depend on the choice of market variables of convenience.
In what follows we propose an approach to defining arbitrarily large correlation stress tests. The approach uses the geometric structure of the space of correlation stress tests, which turns out to form a geodesically complete sub-manifold of the Fisher–Rao manifold of multivariate normal distributions. As a result, there is a natural measure of the size of the stress, as well as a measure of its plausibility. The general conclusions are independent of the selection of particular coordinates on the Fisher–Rao manifold, and thus are geometrically invariant. In particular, they describe stress tests that are independent of the holdings of a portfolio or the representation of the risk of these holdings when expressed using different market factors. This is useful when one aims to apply stress tests consistently across different portfolios. The proposed approach ends up having three useful features. First, it is exhaustive, in the sense that it allows one to explore all correlation stress tests. Second, it is universal, in the sense that the same stress tests can be applied to any covariance matrix. Third, it is quantifiable, in the sense that it provides a natural measure of plausibility of a stress test.
This note is organised as follows: the next section introduces certain well-known features of the Fisher–Rao manifold. Section 3 discusses the meaning of covariance and correlations stress tests and Section 4 discusses the interpretation of correlation stress tests as geodesics in the Fisher–Rao manifold. Section 5 proposes an explicit parametrisation of the correlation stress tests which is the exponential map of the tangent space at the un-stressed covariance. Section 6 discusses the plausibility of the stress tests, while sections 7 through 9 discuss the stress tests that result from specific choices of “directions” in which the correlation matrix is stressed. In Section 10 we show that a correlation stress test which affects all correlations by the same amount results in the same lowest-order functional form as the estimation bias of eigenvalues of the covariance matrix due to finite amount of data. Section 11 reflects on certain features of the stress tests defined by a Lax-pair evolution that would make them less suitable as a model for correlation stress test. Section 12 illustrates the main concepts using a stylised example of bonds, equities and commodities. The results are summarised in Section 13.
2. The Fisher–Rao manifold
The Fisher–Rao metric for a family of probability distributions $p(X;\theta_{i})$ on an $n$-dimensional space of random variables $X$ parametrised by $\theta_{i}$, $i=1,\dots,k$ is given by [10, 11]:
[…]
It is, under fairly general assumptions, the unique metric that does not depend on either the selection of the underlying variables and or on the representation of the parameters of the distributions. The corresponding length is known as the Rao length. Its line element, with a slight abuse of notation, can be expressed as
[…]
More specifically, consider the family of multivariate Gaussian distributions with vector of means $\mu$ and covariance $\Sigma$, and assume that both $\mu$ and $\Sigma$ are differentiable functions of some arbitrary parameters $\theta_{i}$, $i=1,\dots,k$ The Fisher–Rao metric reduces to [12, 13]:
$\begin{split}g_{ij}(\theta)=&\frac{\partial\mu^{T}}{\partial\theta_{i}}\Sigma^{-1}\frac{\partial\mu}{\partial\theta_{j}}+\\ &\frac{1}{2}\operatorname{Tr}\left(\Sigma^{-1}\frac{\partial\Sigma}{\partial\theta_{i}}\Sigma^{-1}\frac{\partial\Sigma}{\partial\theta_{j}}\right).\end{split}$ (4)
In what follows we will consider the case of $\partial\mu/\partial\theta_{i}=0$ for all $i$, i.e. the submanifold of multivariate normal distributions with constant means. Then, the Fisher–Rao distance between two distributions with covariances $\Sigma_{2}$ and $\Sigma_{1}$ in the $n(n+1)/2$-dimensional totally geodesic submanifold of Gaussian distributions with constant drifts is [12]:
$d^{2}_{\mu}(\Sigma_{1},\Sigma_{2})=\frac{1}{2}\sum_{i=1}^{n}\left[\log{(\lambda_{i})}\right]^{2},$ (5)
where $\lambda_{i}$ are the eigenvalues of $\Sigma_{1}^{-1}\Sigma_{2}$. The unique curve of the shortest Rao length, the geodesic $\gamma(t)$, that connects two sufficiently close matrices $\Sigma_{1}$ at $t=0$ to $\Sigma_{2}$ at $t=1$ lies entirely within the multivariate Gaussian submanifold with constant means of the (full) Fisher–Rao manifold. It can be explicitly constructed as [14, 13]:
$\gamma(t)=\Sigma_{1}^{\frac{1}{2}}\exp\left[t\log\left(\Sigma_{1}^{-\frac{1}{2}}\Sigma_{2}\Sigma_{1}^{-\frac{1}{2}}\right)\right]\Sigma_{1}^{\frac{1}{2}},$ (6)
where $\Sigma_{1}^{\frac{1}{2}}$ denotes the unique symmetric positive-definite matrix such that $(\Sigma_{1}^{\frac{1}{2}})^{2}=\Sigma_{1}$.
For a recent review of the geometry of probability distributions and some if its applications, see Ref. [15] for the slightly more general case of elliptically symmetric distributions, or Ref. [16] for an overview of applications for covariance matrix estimation and classification using distance between statistical models and samples.
3. Covariance and correlation stress tests
A covariance between risk factors can be thought of as a symmetric positive definite bilinear form on the space of risk factors. A change of basis in the space of risk factors will result in a change of entries in the matrix representing the covariance form, but this will not change the covariance form itself.
However, a change of entries of the covariance matrix which does not result from a change basis will correspond to a different covariance form. Any change to the covariance form can be thought of as a covariance stress test. We will assume that the stress tests depend on a parameter $t$ that can be interpreted as corresponding to the size of the stress.
We will define a covariance stress test as a one-parameter family of functions that map the space of symmetric positive definite bilinear forms in $\mathbb{R}^{n}$ into itself. A correlation stress test will be a covariance stress that does not change the determinant of the covariance matrix along each path within the family. Thus, we define a correlation stress test as a covariance stress that preserves the generalized variance of Wilks [17].
For convenience, $\Sigma(0)$ will refer to the unstressed, or original, covariance matrix, while $\Sigma(t)$ will be the stressed covariance matrix.
One could object to calling a so-defined stress a correlation stress test, and require instead that all the diagonal elements of the covariance matrix remain unchanged. The diagonal elements of the covariance matrix are defined with respect to a specific basis, so this would not be a basis-independent definition. Moreover, for any two symmetric positive definite matrices with the same determinant, there exists a basis [18] in which they have the same diagonal elements. Thus, a covariance stress test which does not change the determinant of the covariance matrix can always be interpreted as keeping, in some basis, all of the diagonal elements the same and just changing the correlations, which is consistent with the intuitive understanding of stressing correlations. In conclusion, requiring no change of the determinant results in defining a correlation stress test as a covariance stress test that, in some basis, only changes only the correlations between risk factors while keeping all variances the same.
A geometric interpretation provides some further guidance: the determinant of the covariance matrix is, for a multivariate normal distribution, the volume of an ellipsoid of constant probability density enveloping a given probability mass. The directions of the semi-axes of this ellipsoid represent the uncorrelated directions (principal components), while the length of a semi-axis represents the standard deviation of the principal component. A change in the covariance matrix that preserves the determinant would thus preserve the probability mass within it, or its generalised variance, while potentially changing the directions, and possibly also lengths, of the semi-axes. This seems to be the closest analogue to the simple intuitive notion that can be formulated without reference to a specific choice of variables used to represent the probability distribution.
Equivalently, we could define a correlation stress test as a change in the covariance matrix that does not change the entropy $S=\frac{n}{2}+\frac{n}{2}\ln{2\pi}+\frac{1}{2}\ln{\det\Sigma}$ of the Gaussian distribution with covariance $\Sigma$. As our correlation stress tests preserve entropy, they could also be called adiabatic stress tests in analogy with thermodynamics, where the adiabatic processes do not change entropy. This definition is somewhat more general than the one that refers explicitly to the covariance matrix, since entropy can be defined for any probability distribution, $S=-E\left(\log{p}\right)$.
An isotropic stress which scales all the elements of the covariance matrix by the same positive number obviously does not change any correlations. Given two covariance matrices, one can always rescale one of them so that its determinant is equal to the determinant of the other. Thus, any covariance stress can be expressed as a correlation stress defined above and scaling of all entries of the covariance matrix by the same amount.
The following stylised example is perhaps useful in providing some further intuition. The example shows that what is clearly a correlation stress test in one basis ends up changing the variances in another, equally natural, basis. Consider a portfolio that consists of equities and bonds. Assume the volatility of equities is 12% and the volatility of price of bonds is 6%, and the bond–equity return correlation is 0. It is natural to consider returns of equities and bonds as the risk factors, which corresponds to the bond–equity set of market factors. But or course this is not a unique set of possible market factors for this portfolio. Now, consider another set of variables that could be used: a combination of returns of a balanced portfolio of 60% of equities and 40% of bonds, and the equity-bond spread, i.e., a factor which is 100% equity returns less 100% bond returns. Let’s call it a balanced–spread set of market factors. As described, this new set of market factors has been derived from the original set by a change of variables. The original covariance matrix in the bond–equity basis is:
$\Sigma=0.0001\begin{pmatrix}144&0\\ 0&36\\ \end{pmatrix}.$ (7)
In the balanced–spread basis the covariance matrix is
$\Sigma=0.0001\begin{pmatrix}57.6&72\\ 72&180\\ \end{pmatrix}.$ (8)
Now suppose the correlation between bonds and equities is stressed to $0.1$. The respective covariance matrices become
$\Sigma=0.0001\begin{pmatrix}144&7.2\\ 7.2&36\\ \end{pmatrix}$ (9)
in the bond–equity basis, and
$\Sigma=0.0001\begin{pmatrix}61.056&70.56\\ 70.56&165.60\\ \end{pmatrix}$ (10)
in the balanced–spread basis. What appears to be a correlation stress test in the bond–equity basis does not seem to appear as a correlation stress in the balanced–spread basis since the diagonal elements in the latter basis have changed. Additionally, the determinant of the covariance matrix has changed from $5.184\times 10^{-5}$ to $5.132\times 10^{-5}$ (to three decimal places), and thus the probability mass enveloped by the ellipsoid of a constant probability density, as discussed above, has also changed. This example shows that if one aims for correlation stress tests that are more universal and do not depend on such arbitrary choices of basis of risk factors as in the above example, the intuitive approach needs to be refined.
As a side comment, one may observe that given two symmetric positive definite matrices $\Sigma_{1}$ and $\Sigma_{2}$ there always exists a basis such that the entries of $\Sigma_{2}$ in this new basis are the same as the entries of $\Sigma_{1}$ in the original basis. This can be shown by applying the change of basis $V=\Sigma_{2}^{-\frac{1}{2}}\Sigma_{1}^{\frac{1}{2}}$, since
$V^{T}\Sigma_{2}V=\left(\Sigma_{2}^{-\frac{1}{2}}\Sigma_{1}^{\frac{1}{2}}\right)^{T}\Sigma_{2}\left(\Sigma_{2}^{-\frac{1}{2}}\Sigma_{1}^{\frac{1}{2}}\right)=\Sigma_{1}.$ (11)
Thus, any stress test of the covariance matrix, naturally interpreted as an “active” change of the covariance matrix, can always be represented as a “passive” change of the basis in the vector space of the market risk factors. Such a change in general involves stretching the basis vectors, as well as rotating them. It cannot be represented by a purely orthogonal transformation and isotropic scaling except for the trivial case in which the covariance matrices commute, $\left[\Sigma_{1},\Sigma_{2}\right]=0$. As a result, at a fixed point in the Fisher–Rao manifold, the space of infinitesimal covariance stress tests is isomorphic with the group $GL(n,\mathbb{R})$, while the space of correlation stress tests is isomorphic with the group $SL(n,\mathbb{R})$. In a sense, any covariance stress test can be thought of as “some market factors just start behaving like some other market factors”, for example bonds beginning to move more like equities.
4. Correlation stress tests as geodesics in the Fisher–Rao manifold
Since
$\det\left(\exp{A}\right)=\exp\left(\operatorname{Tr}A\right)$ (12)
for any square matrix $A$, and
$\det B=\exp{\left(\operatorname{Tr}\log{B}\right)}$ (13)
for an invertible positive-definite matrix $B$, it follows from Eq. 6 that the determinant of a covariance matrix along a geodesic path is a simple function of the parameter $t$,
$\det\gamma(t)=$ $\det\left[\Sigma_{1}^{\frac{1}{2}}\exp\left[t\log\left(\Sigma_{1}^{-\frac{1}{2}}\Sigma_{2}\Sigma_{1}^{-\frac{1}{2}}\right)\right]\Sigma_{1}^{\frac{1}{2}}\right]$
$=$ $\det\left[\exp\left[t\log\left(\Sigma_{1}^{-\frac{1}{2}}\Sigma_{2}\Sigma_{1}^{-\frac{1}{2}}\right)\right]\right]\det\Sigma_{1}$
$=$ $\exp{\left[t\operatorname{Tr}\left(\log\left(\Sigma_{1}^{-\frac{1}{2}}\Sigma_{2}\Sigma_{1}^{-\frac{1}{2}}\right)\right)\right]}\det\Sigma_{1}$
$=$ $\left(\exp{\left[\operatorname{Tr}\left(\log\left(\Sigma_{1}^{-\frac{1}{2}}\Sigma_{2}\Sigma_{1}^{-\frac{1}{2}}\right)\right)\right]}\right)^{t}\det\Sigma_{1}$
$=$ $\left(\det\left(\Sigma_{1}^{-\frac{1}{2}}\Sigma_{2}\Sigma_{1}^{-\frac{1}{2}}\right)\right)^{t}\det\Sigma_{1}$
$=$ $\det\left(\Sigma_{1}^{-1}\right)^{t}\left(\det\left(\Sigma_{2}\right)\right)^{t}\det\Sigma_{1}$
$=$ $\left(\det\Sigma_{1}\right)^{1-t}\left(\det\Sigma_{2}\right)^{t}.$
In particular, a geodesic joining two covariance matrices whose determinant is the same consists of matrices of the same determinant.
$\det\gamma(t)=\det\Sigma_{1}=\det\Sigma_{2}.$ (14)
Select a point $\Sigma_{1}$ in the Fisher–Rao manifold of Gaussian distributions corresponding to all covariance matrices. The sub-manifold passing through $\Sigma_{1}$ that consists of all matrices $\Sigma$ with the same determinant, $\det\Sigma=\det\Sigma_{1}$, is thus spanned by the geodesics emanating from $\Sigma_{1}$ that lie entirely within that submanifold, and so define an exponential map within that submanifold. As a result of the geodesics always staying within this submanifold, the extrinsic curvature of the submanifold of covariance matrices with constant determinant, as embedded in the Fisher–Rao manifold, vanishes [19].
In the next section, we will show that the exponential map can be extended to arbitrarily long geodesics, and thus defines arbitrarily large correlation stress tests.
5. Parametrisations of the sub-manifold of the correlation stress tests
Skovgaard’s [20] Theorem 6.7 shows how a geodesic can be constructed by diagonalisation with respect to the original covariance matrix, but this original formulation may not have an apparent intuitive link to the correlation stress tests. It is perhaps more natural to consider, as in Ref. [14], a one-parameter family of covariance matrices $\Sigma(t)$ which are a result of stressing $\Sigma=\Sigma(0)$ to $\Sigma(t)$
$\Sigma(t)=\Sigma_{X}(t)=\Sigma^{\frac{1}{2}}\exp{(Xt)}\Sigma^{\frac{1}{2}},$ (15)
where the subscript in $\Sigma_{X}(t)$ indicates the explicit dependence of this family on a symmetric traceless matrix $X$, and the parameter $t$ corresponds to the amount of stress. Thus, $\Sigma(t)$ is symmetric and positive definite. Also, since
$\det(\exp(Xt))=\exp(\operatorname{Tr}(Xt))=1,$ (16)
$\det\Sigma(t)=\det\Sigma(0)$, so Eq. 15 indeed defines a correlation stress and a geodesic in the Fisher–Rao manifold.
This can be explicitly confirmed substituting Eq. 15 into Eq. 6 and setting $\Sigma_{1}=\Sigma(0)$ and $\Sigma_{2}=\Sigma(1)$ to show [14] that
$\gamma(t)$ $=\Sigma_{1}^{\frac{1}{2}}\exp{\left[t\log{\left(\Sigma_{1}^{-\frac{1}{2}}\Sigma_{1}^{\frac{1}{2}}\exp{(X)}\Sigma_{1}^{\frac{1}{2}}\Sigma_{1}^{-\frac{1}{2}}\right)}\right]}\Sigma^{\frac{1}{2}}$
$=\Sigma^{\frac{1}{2}}\exp\left[t\log\exp{(X)}\right]\Sigma^{\frac{1}{2}}$
$=\Sigma^{\frac{1}{2}}\exp{(Xt)}\Sigma^{\frac{1}{2}}$
$=\Sigma(t).$ (17)
Thus Eq. 15 is indeed a geodesic parametrised by $t$. Alternatively, this can be confirmed by applying Theorem 6.1 of Ref. [20] to $\Sigma(t)$ defined in Eq. 15. For a choice of basis in the space of matrices $X$ that span the space of all traceless symmetric matrices of order $n$, Eq. 15 defines explicitly the exponential map in the sub-manifold of correlation stress tests.
Eq. 15 is formally equivalent to Eq. 6, but rather than depending explicitly on the end points of the geodesic, it is explicitly parametrised by the starting point and an arbitrary trace-free symmetric matrix which corresponds to the tangent space to the Fisher–Rao manifold at the point $\Sigma(0)$. In addition to the path defined in Eq. 15, consider another path parametrised by a trace-free symmetric matrix $Y$,
$\Sigma_{Y}(t)=\Sigma^{\frac{1}{2}}\exp{(Yt)}\Sigma^{\frac{1}{2}}.$ (18)
The scalar product at $\Sigma(0)$ defined by Eq. 4 reduces to
$g(X,Y)=$ $\frac{1}{2}\operatorname{Tr}\left(\Sigma^{-1}\frac{\partial\Sigma_{X}}{\partial t}\Sigma^{-1}\frac{\partial\Sigma_{Y}}{\partial t}\right)$ (19)
$=$ $\frac{1}{2}\operatorname{Tr}(XY),$ (20)
i.e. the Frobenius norm on the space of the traceless matrices $X$ that define tangent space at $\Sigma$. Equation 15 can of course be also written as
$\Sigma(t)=\Gamma^{T}(t)\Gamma(t)$ (21)
with $\Gamma(t)=\exp{\left(\frac{1}{2}Xt\right)}\Sigma^{\frac{1}{2}}$.
As $\Sigma^{\frac{1}{2}}$ is the matrix of change of basis into one in which the covariance matrix is the identity matrix, one can interpret the above parametrisation as referring, implicitly, to changing the correlations away from the identity matrix, which is its generic form of correlations when expressed in the basis of principal components.
Various choices of $X$ can be made. A specific selection of the entries of $X$ will define a stress test with respect to a specific choice of basis in the space of the random variables. A practitioner may have a portfolio of special concern, whose exposures may be known in a defined basis. Or perhaps, considering a narrative of what may affect the markets, they may be concerned with the behaviour of a number of portfolios, some of which may be affected in different ways. In the following sections, we will consider a measure of size of a stress test — its plausibility, which allows one to have an a priori view on what size stress tests need to be considered, without specifying the portfolio and its market factors themselves.
6. Plausibility of correlation stress tests
The line element defined by Eq. 3 is the expected value of the change of log-probability. Thus, the geodesic distance $d$ between two covariance matrices may be thought of as the integral of the infinitesimal differences between their log-probabilities, so Eq. 5 defines a natural notion of the plausibility of a stress test that morphs $\Sigma_{1}$ into $\Sigma_{2}$:
$P\left(\Sigma_{1}\rightarrow\Sigma_{2}\right)=\exp{(-d)}$ (22)
For a specific correlation stress test path defined in Eq. 15, this distance, Eq. 5, can be calculated using the eigenvalues $\lambda_{i}$ of $\Sigma^{-1/2}\exp{(Xt)}\Sigma^{1/2}$. Let $x_{i}$ be an eigenvalue corresponding to an eigenvector $v_{i}$ of the above matrix, so the eigenvalue equation is
$\Sigma^{-1/2}\exp{(Xt)}\Sigma^{1/2}v_{i}=x_{i}v_{i}.$ (23)
Multiplying both sides by $\Sigma^{1/2}$ and setting $\Sigma^{1/2}v_{i}=u_{i}$ shows that the eigenvalues $x_{i}$ satisfy
$\exp{(Xt)}u_{i}=x_{i}u_{i}.$ (24)
Thus $x_{i}$ is an eigenvalue of $\exp(Xt)$, or $\log(x_{i})$ is an eigenvalue of $X$ multiplied by $t$. Then Eq. 5 implies
$d^{2}\left(\Sigma(0),\Sigma(t)\right)=\frac{t^{2}}{2}\sum_{i=1}^{n}x_{i}^{2}$ (25)
with $x_{i}$ being the eigenvalues of $X$. Since the values of $d^{2}$ can be arbitrarily large for some $t\in\mathbb{R}$, the submanifold of correlation stress tests is geodesically complete.
Using Eq. 22, the plausibility of a correlation stress test defined by stressing $\Sigma(0)$ to $\Sigma(t)$ as defined by Eq. 15 is
$P(X,t)=\exp{\left(-\sqrt{\frac{t^{2}}{2}\sum_{i=1}^{n}x_{i}^{2}}\right)}.$ (26)
More plausible stress tests would perhaps be higher concern to a risk manager, although it is not apparent at what absolute level of of plausibility such concerns should become acute. Specifying such a level, between zero and one, is akin to explicitly specifying a level of risk tolerance to stress tests. In addition, Eq. 26 can be used to assess relative plausibility of one stress test compared to another. For example, a stress test that is tenfold, or hundredfold, as plausible as another stress test, may perhaps be considered as more relevant. Note that the plausibility of such stress tests does not depend on the initial covariance matrix as $\Sigma$ does not enter Eq. 26. Thus, the plausibility of a correlation stress test is in a sense universal: different covariance matrices can be subject to the same correlation stress test of a given plausibility.
7. Stressing a single pair of correlations
If one selects a specific basis of the random variables, one can think of stressing a correlation of two variables indexed by $a$ and $b$, $a\neq b$, as defined by by Eq. 15 with $X_{ij}=\delta^{a}_{(i}\delta^{b}_{j)}$ where the brackets denote symmetrisation with respect to the respective indices. Without loss of generality, one can assume that $a=1$ and $b=2$, and represent $X$ in a block-diagonal form where the entries with $X_{i,j}=0$ for $i,j>2$ except for a block where $i\leq 2$ and $j\leq 2$. Then
$X=\begin{pmatrix}0,&1,&0,&\cdots&,&0\\ 1,&0,&0,&\cdots&,&0\\ 0,&0,&0,&\cdots&,&0\\ &&\cdots&&&\\ 0,&0,&0,&\cdots&,&0\\ \end{pmatrix},$ (27)
and
$X^{2}=\begin{pmatrix}1,&0,&0,&\cdots&,&0\\ 0,&1,&0,&\cdots&,&0\\ 0,&0,&0,&\cdots&,&0\\ &&\cdots&&&\\ 0,&0,&0,&\cdots&,&0\\ \end{pmatrix}.$ (28)
As a consequence, $\exp{(tX)}$ is also block-diagonal with the same structure,
$e^{tX}=\begin{pmatrix}\cosh{t},&\sinh{t},&0,&\cdots&,&0\\ \sinh{t},&\cosh{t}&0,&\cdots&,&0\\ 0,&0,&1,&\cdots&,&0\\ &&\cdots&&&\\ 0,&0,&0,&\cdots&,&1\\ \end{pmatrix}.$ (29)
If, in addition, in that specific basis the covariance matrix $\Sigma$ is diagonal, $\Sigma=\operatorname{diag}(\sigma_{i}^{2})$, then the stressed covariance matrix $\Sigma(t)$ is itself block-diagonal
$\Sigma(t)=\begin{pmatrix}\sigma_{1}^{2}\cosh{t},&\sigma_{1}\sigma_{2}\sinh{t},&0,&\cdots&,&0\\ \sigma_{1}\sigma_{2}\sinh{t},&\sigma_{2}^{2}\cosh{t}&0,&\cdots&,&0\\ 0,&0,&\sigma_{3}^{2},&\cdots&,&0\\ &&\cdots&&&\\ 0,&0,&0,&\cdots&,&\sigma_{n}^{2}\\ \end{pmatrix}.$ (30)
The entries $\sigma_{j}^{2}$ for $j>2$ are unaffected by the stress test. While the infinitesimal stress test $X$ only impacts the two entries corresponding to the correlations of the two variables, in a finite stress test, all covariances in the top-left block, including the diagonal elements, are affected by this correlation stress, although of course the determinant of the covariance matrix does not change. If the stress test is with respect to a basis in which the covariance matrix is diagonal, only the block corresponding to the variables whose correlation is stressed is affected, while the remaining covariances stay the same, which one would intuitively expect. This is not the case if $\Sigma^{\frac{1}{2}}$ is not diagonal, as all correlations are effectively mixed and interrelated in such stress tests.
Note that in the above stress test, the diagonal elements are unchanged up to first order of $t$, thus the variances and standard deviations are not affected up to order $O(t)$ by the infinitesimal version of this stress test.
8. Stressing along the diagonal
Of course a traceless matrix $X$ needs not have all zero diagonal elements. Without loss of generality, a convenient choice of a basis of the stress tests within the diagonal can be parametrised by setting $X_{11}=1$ and $X_{22}=-1$, with all other entries being 00:
$X=\begin{pmatrix}1,&0,&0,&\cdots&,&0\\ 0,&-1,&0,&\cdots&,&0\\ 0,&0,&0,&\cdots&,&0\\ &&\cdots&&&\\ 0,&0,&0,&\cdots&,&0\\ \end{pmatrix}.$ (31)
Thus
$X^{2}=\begin{pmatrix}1,&0,&0,&\cdots&,&0\\ 0,&1,&0,&\cdots&,&0\\ 0,&0,&0,&\cdots&,&0\\ &&\cdots&&&\\ 0,&0,&0,&\cdots&,&0\\ \end{pmatrix}.$ (32)
As a consequence, $\exp{(tX)}$ is also block-diagonal,
$e^{tX}=\begin{pmatrix}\exp{(t)},&0,&0,&\cdots&,&0\\ 0,&\exp{(-t)},&0,&\cdots&,&0\\ 0,&0,&1,&\cdots&,&0\\ &&\cdots&&&\\ 0,&0,&0,&\cdots&,&1\\ \end{pmatrix}.$ (33)
If such a stress is applied to a covariance matrix $\Sigma$ that is diagonal, $\Sigma=\operatorname{diag}(\sigma_{i}^{2})$, the stressed covariance matrix $\Sigma(t)$ is itself block-diagonal
$\Sigma(t)=\begin{pmatrix}\sigma_{1}^{2}\exp{(t)},&0,&0,&\cdots&,&0\\ 0,&\sigma_{2}^{2}\exp{(-t)}&0,&\cdots&,&0\\ 0,&0,&\sigma_{3}^{2},&\cdots&,&0\\ &&\cdots&&&\\ 0,&0,&0,&\cdots&,&\sigma_{n}^{2}\\ \end{pmatrix}.$ (34)
The stress test described above may not be intuitively perceived as a correlation stress. However, together with the stress tests described in the previous section, the matrices $X$ may be used to form a basis of the tangent space at any covariance matrix $\Sigma(0)$. Thus, the examples in these two sections describe a basis of the most general correlation stress tests.
9. Stressing all correlations of a single variable
In a selected basis of the random variables, one can think of stressing all correlations of a single factor with all other factors by an equal amount, with all other correlations unchanged. Without loss of generality, assume that this variable is the first variable, so that $X$ can be represented as:
$X=\begin{pmatrix}0,&1,&1,&\cdots&,&1\\ 1,&0,&0,&\cdots&,&0\\ 1,&0,&0,&\cdots&,&0\\ &&\cdots&&&\\ 1,&0,&0,&\cdots&,&0\\ \end{pmatrix},$ (35)
so
$X^{2}=\begin{pmatrix}n-1,&0,&0,&\cdots&,&0\\ 0,&1,&1,&\cdots&,&1\\ 0,&1,&1,&\cdots&,&1\\ &&\cdots&&&\\ 0,&1,&1,&\cdots&,&1\\ \end{pmatrix}.$ (36)
Let $Z=X^{2}$. Then
[…]
The power series expansion of $\exp{(tX)}$ can then be expressed as a linear combination of X and Z,
$\exp{(tX)}=$ $I+\frac{1}{\sqrt{n-1}}X\sinh{(\sqrt{n-1}t)}$ (38)
$+\frac{1}{n-1}Z\left(\cosh{(\sqrt{n-1}t)}-1\right)$ (39)
where $I$ is the $n\times n$ identity matrix.
10. Equally stressing all correlations
In the previous sections we have discussed several stress tests resulting from specific choices of the matrix $X$ generating the paths given by Eq. 15. In this section we will consider a hollow matrix $X$ whose non-diagonal entries are all equal. Lawley [21] discussed the the bias of estimates of eigenvalues of the covariance matrix due to finite data sample size. He showed that the estimates $\hat{\lambda}_{r}$ of the largest eigenvalues $\lambda_{r}$ of the covariance matrix $\Sigma$ of $p$ random variables are related to the the eigenvalues $l_{r}$ of the sample covariance matrix with data sample of size $m$ by
$\hat{\lambda}_{r}=$ $l_{r}\left\{1-\frac{1}{m}\sum^{k}_{i=1;i\neq r}\frac{l_{i}}{l_{r}-l_{i}}-\frac{p-k}{m}\frac{\lambda}{l_{r}-\lambda}\right\}$
$+O\left(\frac{1}{m^{2}}\right),$ (40)
where the remaining $p-k$ smaller eigenvalues are all assumed to be equal to $\lambda$. This equation has a notable feature: if $p=k$, the trace of the sample covariance matrix is equal to the trace of the population covariance matrix, up to order of $O\left(m^{-2}\right)$.
This observation was used in [22] to propose a stress test defined by formally varying $m$ in Eq. 10,
$\lambda_{r}=l_{r}\left[1-s\sum^{p}_{i=1;i\neq r}\frac{l_{i}}{l_{r}-l_{i}}\right],$ (41)
where the parameter $s$ defines the size of the stress test.
The stress above, obtained by varying $s$, does not change the trace of the covariance matrix, hence up to the first order of $s$ it does not change its determinant. Thus it can be interpreted as an infinitesimal correlation stress test. However, it can be applied only when $s$ is small, and thus allows only for small correlation stress tests. Large stress tests, corresponding to a large value of $s$, could result in non-positive $\lambda_{r}$. In addition, this stress has only one parameter, and thus is cannot capture the full richness of the possible changes of the covariance matrix such as defined by Eq. 15. It is however an infinitesimal version of a specific stress test of Eq. 15, namely one with all off-diagonal elements of $X$ being equal. This can be shown as follows: set
$X=\begin{pmatrix}0,&1,&1,&\cdots&,&1\\ 1,&0,&1,&\cdots&,&1\\ 1,&1,&0,&\cdots&,&1\\ &&\cdots&&&\\ 1,&1,&1,&\cdots&,&0\\ \end{pmatrix},$ (42)
in the basis of the eigenvectors of $\Sigma(0)$, so $\Sigma(0)=\operatorname{diag}(\sigma_{i}^{2})$. This corresponds to setting $l_{r}=\sigma_{r}^{2}$ in Eq. 41. Consider the first two terms in the power series expansion in $t$ of Eq. 15:
$\Sigma(t)=\Sigma^{\frac{1}{2}}\left(I+Xt\right)\Sigma^{\frac{1}{2}}.$ (43)
By using the power series expansion in $t$ of the eigenvalues of $\Sigma(t)$ we will show that they satisfy Eq. 41, with $s=2t$, up to order of $O(t^{2})$.
Ref. [23] discusses the well-known feature of the impact of perturbations on the spectrum of symmetric positive-definite matrices: the eigenvalues $\lambda_{i}$ of a one-parameter family of matrices $A(t)$ behave as if they were subject to a mutually repulsive force:
$\dot{\lambda}_{i}=u_{i}^{T}\dot{A}u_{i}$ (44)
and
$\ddot{\lambda}_{i}=u_{i}^{T}\ddot{A}u_{i}+2\sum_{j\neq i}{\frac{|u_{j}^{T}\dot{A}u_{i}|^{2}}{\lambda_{i}-\lambda_{j}}},$ (45)
where the dot denotes differentiation with respect to the parameter $t$, two dots denote the second derivative, and $u_{i}$ are the normalised eigenvectors of $A(0)$.
Set $A(t)=\Sigma(t)$, and let $u_{i}$ be the normalised eigenvectors of $\Sigma(0)$. The first derivative of $\Sigma(t)$ given by Eq. 43 is
$\dot{\Sigma}(0)=\Sigma^{\frac{1}{2}}X\Sigma^{\frac{1}{2}},$ (46)
where $\Sigma^{\frac{1}{2}}=\operatorname{diag}(\sigma_{i})$. All diagonal entries of the matrix $X$, and hence of $\dot{\Sigma}(0)$, are zero, all off-diagonal entries of $\dot{\Sigma}(0)$ are $\sigma_{j}\sigma_{k}$, and all components of the vectors $u_{i}$ are zero except for the $i$-th component which is $1$. As a result $u_{i}^{T}\dot{\Sigma}(0)u_{i}=0$, while $u_{j}^{T}\dot{\Sigma}(0)u_{k}=\sigma_{j}\sigma_{k}$ for $j\neq k$. Separately, the second derivative of $\Sigma(t)$ given by Eq. 43 is zero. These observations have the following two implications.
First, as a result of Eq. 44, the first derivative of the eigenvalues of $\Sigma(t)$ defined by Eq. 43 vanish at $t=0$. Hence, there is no first-order term in the power-series expansion of the eigenvalues of $\Sigma(t)$ around $t=0$.
Second, since $\ddot{\Sigma}(t)$ of Eq. 43 vanishes, Eq. 45 shows that the second derivative of the eigenvalues of $\Sigma(t)$ defined by Eq. 43 is
$\ddot{\lambda}_{i}=2\sum_{j\neq i}\frac{\sigma_{i}^{2}\sigma_{j}^{2}}{\lambda_{i}-\lambda_{j}}=2\sum_{j\neq i}\frac{\lambda_{i}\lambda_{j}}{\lambda_{i}-\lambda_{j}}.$ (47)
As a result, the first two terms in the power series expansion around $t=0$ of the eigenvalues of $\Sigma(t)$ have the same functional form as Eq. 41, with $s=\sqrt{t}$.
Interestingly, the original article of Lawley [21] considered the bias of eigenvalues due to finite amount of data. One can interpret the above result as implying that increasing (or decreasing) the amount of data may result in what is effectively a correlation stress. Eq. 15 with $X$ of Eq. 42 allows such stress test to be extended to an arbitrarily large size, or equivalently to arbitrarily low plausibility.
This type of correlation stress may also be useful for stress-testing of the valuation of derivatives with a relatively large number of similar underlying market factors, for example a swap on volatility or variance of a basket or an option on a basket. For such baskets, the covariance matrix is of course likely to be non-diagonal, but the stress test defined by $X$ in Eq. 42 can still be applied.
It may be worth noting that, for $X$ defined in Eq. 42, there is a relatively simple expression for $\exp{\left(tX\right)}$. Adding the identity matrix $I$ to $X$ results in a matrix of ones $J$, which has a rank of one and has only two eigenvalues, $n=\dim(X)$ of multiplicity of one, and zero as the remaining eigenvalues. By inspecting its eigenvectors, the matrix exponential of $tJ$ can be demonstrated to be
$\exp{\left(tJ\right)}=I+\frac{e^{nt}-1}{n}J,$ (48)
so
$\exp{\left(tX\right)}=\exp{\left(Jt-It\right)}$ (49)
$=e^{-t}\left(I+\frac{e^{nt}-1}{n}J\right).$ (50)
11. Lax pair stress tests
Instead of the approach to correlation stress tests discussed in Section 3, one could require a different, perhaps more stringent, definition of a correlation stress test — namely that the eigenvalues of the covariance matrix do not change under such a stress test. Suppose, for simplicity, that the eigenvalues are distinct, so that the spectrum of $\Sigma(0)$ is non-degenerate. The equality of eigenvalues along a path $\Sigma(t)$ would then mean that both $\Sigma(t)$ and $\Sigma(0)$, when diagonalised, are equal:
$\left(V(t)\right)^{T}\Sigma(t)V(t)=\left(V(0)\right)^{T}\Sigma(0)V(0)$ (51)
for a non-singular orthogonal family of matrices $V(t)$. One can take the derivative of Eq. 51 with respect to $t$, note that $\left(V(t)\right)^{T}=\left(V(t)\right)^{-1}$ because $\Sigma(t)$ remains symmetric, and use
$\frac{dV^{T}}{dt}=\frac{dV^{-1}}{dt}=-V^{-1}\frac{dV}{dt}V^{-1}.$ (52)
This shows that $Y(t)=\frac{dV^{T}}{dt}V^{-1}$ and $\Sigma(t)$ are a Lax pair,
$\frac{d}{dt}\Sigma(t)=[Y(t),\Sigma(t)],$ (53)
where the square brackets denote the commutator of matrices, $[A,B]=AB-BA$.
Eq. 53 is a first order differential equation that defines a path connecting $\Sigma(0)$ with $\Sigma(t)$. Instead of considering, as in Eq. 15, the geodesic paths, one could then perhaps consider stress tests defined by the Lax pair, Eq. 53.
We will now show that in most cases these two equations cannot describe the same path. We then will argue that this is a significant drawback of the Lax pair stress tests.
Suppose that the eigenvalues $\lambda_{i}(t)$ of $\Sigma(t)$ do not change in the neighbourhood of $t=0$. Select a basis of eigenvectors $u_{i}$ of $\Sigma(0)$. In this basis $\Sigma=\Sigma(0)=\operatorname{diag}(\lambda_{i})=\operatorname{diag}(\sigma_{i}^{2})$ and
$\dot{\Sigma}(0)=\Sigma^{\frac{1}{2}}X\Sigma^{\frac{1}{2}}.$ (54)
Let $X_{ij}$ be the entries of the matrix $X$.
Eq. 44 implies that
$\dot{\lambda}_{i}=u_{i}^{T}\dot{\Sigma(t)}u_{i}=u_{i}^{T}\Sigma^{\frac{1}{2}}X\Sigma^{\frac{1}{2}}u_{i}=X_{ii}^{2}\sigma^{2}_{i}.$ (55)
Vanishing of $\dot{\lambda}_{i}$ then implies that the diagonal elements of $X$ vanish, $X_{ii}=0$. Suppose that the spectrum of $\Sigma(0)$ is non-degenerate, so that no two eigenvalues $\lambda_{i}$ are equal. Select the eigenvalues in a decreasing order, so that $\lambda_{1}$ is the largest one. Then Eq. 45 with
$\ddot{\Sigma}(0)=\Sigma^{\frac{1}{2}}X^{2}\Sigma^{\frac{1}{2}}$ (56)
implies that
$\ddot{\lambda}_{i}=$ $u_{i}^{T}\ddot{A}u_{i}+2\sum_{j\neq i}{\frac{|u_{j}^{T}\dot{A}u_{i}|^{2}}{\lambda_{i}-\lambda_{j}}}$
$=$ $\sigma_{i}^{2}\sum_{k}X_{ik}^{2}+2\sum_{k\neq i}\frac{X_{ik}^{2}\sigma_{k}^{2}\sigma_{i}^{2}}{\lambda_{i}-\lambda_{k}}$
$=$ $\sigma_{i}^{2}X_{ii}^{2}+\sum_{k\neq i}\sigma_{i}^{2}X_{ik}^{2}\left(1+2\frac{\sigma_{k}^{2}}{\lambda_{i}-\lambda_{k}}\right)$
$=$ $\sum_{k\neq i}\sigma_{i}^{2}X_{ik}^{2}\left(1+2\frac{\sigma_{k}^{2}}{\lambda_{i}-\lambda_{k}}\right)$
since $X_{ii}=0$.
Start with the largest eigenvalue $\lambda_{1}$. Setting $\ddot{\lambda}_{1}=0$ implies that $X_{1k}=0$ for all $k=1,\dots,n$ since $\lambda_{1}-\lambda_{k}>0$ for all $k$. Since $X$ is symmetric, this also implies $X_{k1}=0$, so $X$ must be a matrix with zeros in the first row and first column. The same argument can then be repeated to show that $X_{2k}=0$ for all $k=1,\dots,n$. Applying this argument recursively shows that setting $\ddot{\lambda_{i}}(0)=0$ for all $i=1,\dots,n$ implies $X=0$. In conclusion, if the spectrum of $\Sigma(0)$ is non-degenerate, the geodesic path that joins it with $\Sigma(1)$ cannot preserve all the eigenvalues.
As a result, the length along the path defined by Eq. 53 will only provide an upper bound on the geodesic length between the stressed and unstressed covariance matrix, or equivalently a lower bound of the plausibility of such a stress test. For practitioners, a lower bound on plausibility may not be very useful as it is difficult to argue in favour of using a stress test which is ”not less plausible” than a certain threshold. Of course, one could always find a geodesic joining two covariance matrices which have the same spectrum, but the intermediate covariance matrices along this geodesic will have a different spectrum.
If the spectrum of $\Sigma(t)$ has repeated eigenvalues, the argument shown above still applies to the directions, in the basis of eigenvectors of $\Sigma(0)$, whose eigenvalues have multiplicities of one, so the entries of $X_{ij}$ are zero in these directions. The subset of such matrices $\Sigma$ is however a lower-dimensional subset of the manifold of correlation stress tests, and thus it cannot correspond to a generic correlation stress test that could be applied to an arbitrary covariance matrix.
12. Example: equities, bonds, and commodities
Suppose one was presented with an explicit stress test, say a given stressed covariance matrix $\Sigma_{2}$. By itself, this would not be a sufficient specification of a covariance stress test, as our approach requires specifying an entire stress path $\Sigma(t)$. However, if a stressed covariance matrix $\Sigma_{2}$ and and non-stressed covariance matrix $\Sigma_{1}$ with respect to some preferred basis of risk factors are specified, one can set the stressed path to be the geodesic joining $\Sigma_{1}$ and $\Sigma_{2}$. This geodesic is the shortest path joining these two matrices, or equivalently, the path which, at each point, points in the direction of the most plausible stress test. If $\det\Sigma_{1}=\det\Sigma_{2}$, the stress test is a correlation stress, and one can calculate
$X=\log{\left(\Sigma_{1}^{-\frac{1}{2}}\Sigma_{2}\Sigma_{1}^{-\frac{1}{2}}\right)}$ (57)
to use in Eq. 15 in order to reconstruct the entire path of stress test.
A practitioner may further be faced with a situation where while $\Sigma_{1}$ is known, the stressed covariance matrix $\Sigma_{2}$ is not fully defined, for example only some of its entries are specified. For example, one can be asked to consider returns of bonds, equities, and commodities, and be asked to stress test the correlation of bonds and equities without any specification of what would happen to the other correlations. In such cases, it may be natural to define the stressed covariance matrix $\Sigma_{2}$ by requiring that it be the most plausible correlation stress test, namely to find the missing stressed entries of $\Sigma_{2}$ by the optimisation problem
$\Sigma_{2}\in\arg\min d^{2}(\Sigma_{1},\Sigma_{2})$ (58)
where $d$ is defined by Eq. 5. This may be particularly useful if only a few stressed correlations are specified.
To illustrate the above, let us return to the stylised example discussed in Section 3, but with an additional market factor, say commodities, with annualised volatility of 25% and uncorrelated to the other two factors. The original covariance matrix in the bond–equity basis is
$\Sigma_{1}=0.0001\begin{pmatrix}[r]144&0&0\\ 0&36&0\\ 0&0&625\\ \end{pmatrix}.$ (59)
Now suppose that the correlation between bonds and equities is, as before, stressed to $0.1$. However, this specification narrative, if taken to mean ceteris paribus, namely that all other entries of the covariance matrix remain unchanged, would change of the determinant of the covariance matrix from $3.2400\times 10^{-6}$ to $3.2076\times 10^{-6}$. On the other hand, mutatis mutandis, if the other entries of the covariance matrix except for the ones for bonds and equities are allowed to vary, this specification of the stress test could mean:
$\Sigma_{2}=0.0001\begin{pmatrix}[r]144&7.2&x\\ 7.2&36&y\\ x&y&z\\ \end{pmatrix},$ (60)
where $x$, $y$, and $z$ are arbitrary, subject to the constraint $\det\Sigma_{1}=\det\Sigma_{2}$ and positive-definiteness of $\Sigma_{2}$.
The values of $x$, $y$, and $z$ that minimise the distance defined in Eq. 58 are 24.76, 23.84, and 649.90, respectively. Thus, the most plausible result of stressing the bond–equity correlation from 0 in $\Sigma_{1}$ to 0.1 is
$\Sigma_{2}=0.0001\begin{pmatrix}[r]144.00&7.20&24.76\\ 7.20&36.00&23.84\\ 24.76&23.84&649.90\\ \end{pmatrix},$ (61)
to two decimal places. As a result of this most-plausible correlation stress test consistent with bond–equity correlation being 0.1, the correlation between equities and commodities becomes 0.08, and the correlation between bonds and commodities becomes 0.16. Note that the volatility of commodities is also affected, but it increases only marginally to 25.49%. This increase of volatility is necessary in order to keep the determinant, or the generalised variance, unchanged, as required for a correlation stress.
The above correlation stress $\Sigma_{2}$ corresponds to the matrix $X$ in Eq. 57 that defines a continuous path of the most plausible stress tests, Eq. 15,
$X=\begin{pmatrix}[r]-0.007613&0.094822&0.074094\\ 0.094822&-0.016781&0.153825\\ 0.074094&0.153825&0.024395\\ \end{pmatrix}.$ (62)
The trace of $X$ may not appear to be zero due to rounding to six decimal places, but it does indeed vanish. The sum of its eigenvalues squared, as in Eq. 25, is 0.077222 to four decimal places. Note that reasonably high numerical accuracy is required for calculations of exponentials of matrices.
In the example above, Eq. 15 can be used to extend the path, defined by $X$, that connects $\Sigma_{1}$ and $\Sigma_{2}$, to an arbitrarily large distance, making it progressively less plausible.
Figure 1: Path of bond–equity-commodity correlation stress tests
Figure 1 shows the behaviour, as a function of $t$, of the Rao distance, plausibility, and the three eigenvalues of the covariance matrix, extended from the original parameter $t=1$ to $t=10$. Note that the distance is just a linear function of $t$ and the plausibility decreases exponentially with $t$. The smallest eigenvalue (that initially corresponds to the volatility of bonds) tends to zero as $t\rightarrow\infty$. While it may not be apparent from the graph, the differences between the eigenvalues initially increase as a function of $t$, as required by the perturbation formulae. Eqs. 44 and 45. The path can be easily extended beyond the range shown in the figure, with numerical accuracy effects becoming non-trivial at $t>125$, with plausibility becoming less than $4\times 10^{-10}$ for $t>120$.
13. Summary
In this note we have discussed certain geometric properties of the Fisher–Rao manifold that can be used to define covariance and correlation stress tests.
These tests are independent of the non-stressed covariance matrix or the portfolio to which they will ultimately be applied. The correlation stress tests can be extended to have arbitrarily large size. We also discussed a natural measure of the absolute and relative plausibility of the stress tests.
We have proposed to use a simple coordinate-independent parametrisation of correlation stress tests, and discussed several specific examples: stressing a pair of correlations, stressing all correlations of a single variable, stressing along the diagonal, and stressing all correlations equally.
The proposed correlation stress tests can be used in calculating stressed risk measures such as ex-ante volatility or parametric value-at-risk, factor sensitivities such as betas, incremental risk measures, etc. For a portfolio of financial assets, such stress tests can be used to assess the values that these risk measures could take reasonably plausible scenarios. For relative-value portfolios, the traditional risk measures sometimes dramatically increase after a market-moving event that results in a meaningful change of the correlations between the positions in the portfolio. The stressed risk measures are potentially less cyclical than the ones traditionally used for limits. Using the stressed risk measures for limits may then allow one to maintain and actively manage such relative-value portfolios, without being forced to reduce the positions just because of realised, but unanticipated, changes of correlations.
The proposed correlation stress tests can also be applied to analyse stressed valuations of financial derivatives whose price depends on implied correlations, such as spread options, binary options contingent on multiple underlying assets, or swaps which depend on the volatility of a basket of assets whose individual constituent assets have active option markets.
I would like to thank Sebastian Armstrong for his comments, in particular for pointing out errors in an earlier version of this note, as well as Juan Antolin-Diaz and the Fulcrum research team for our discussions of this subject. I would also like to thank Witold Biedrzycki for his comments.